OpenStudy (anonymous):

what is the perpendicular line of y=-5x+13

OpenStudy (anonymous):

perpendicular slopes are negative reciprocals so in this case the slope would be 1/5

OpenStudy (anonymous):

but how do i figure out the y intercept or is it going to be the same??

OpenStudy (anonymous):

y=mx+b b is always the y intercept so i this case the y intercept is 13

OpenStudy (anonymous):

so the y intercept doesn't change?

OpenStudy (anonymous):

whenever an equation is in y=mx+b form the b is the y intercept the number changes with every problem though

OpenStudy (anonymous):

i get that b is always the y intercept i mean is the answer going to be y=1/5x+13 or is going to have a different y intercept from 13

OpenStudy (anonymous):

No

OpenStudy (anonymous):

so i can make the y-intercept which ever point i want or is there a way for me to figure it out?

zepdrix (zepdrix):

Hi Sooch :) No the y-intercept will not necessarily be the same. In order to find your new y-intercept, you must be given more information. Usually a problem like this is worded as: "Find the equation of the line which passes through the point ( * , * ) and is perpendicular to y=-5x+13." We weren't given enough information to find your y-intercept.

OpenStudy (anonymous):

my teacher didn't give any points the question just says "the line through the origin that is perpendicular to the graph of y=-5x+13"

zepdrix (zepdrix):

That is a point being given silly! :) The origin is the point (0,0), right?

zepdrix (zepdrix):

So you figured out this much so far.\[\Large\rm y=\frac{1}{5}x+b\]Plug in your (0,0) to find your b value! ^^

OpenStudy (anonymous):

so b is going to equal 0?

zepdrix (zepdrix):

\[\Large\rm 0=\frac{1}{5}\cdot0+b\]Mmm good good!

zepdrix (zepdrix):

\[\Large\rm y=\frac{1}{5}x\]Yay good job \c:/

OpenStudy (anonymous):

thank you so much